# Task-agnostic Pre-training and Task-guided Fine-tuning for Versatile Diffusion Planner

### v1

@@ Appendix A Derivations
A.1 Derivation of policy gradient in Equation  (11)
@@
A.2 Derivation of loss function in Equation (12)
@@ Appendix B The Details of SODP
B.1 Diffusion Policy
@@ B.2 Implementation Details
  • For fine-tuning, we use DDIM (Song et al., 2020) with $10$ sampling steps and $\eta=1$. We fine-tune each task for $1e^{6}$ steps in Meta-World and $3e^{3}$ steps in Adroit. Following DPOK (Fan et al., 2024), we perform $p_{\text{step}}\in\{10,30\}$ gradient steps per episode. We set discount factor $\gamma=1$ for all tasks.
  • We set $N_{\text{init}}\in\{10,20\}$ for approximating target distribution.
@@
       Sample $(\bm{s},\bm{a})\sim D$, diffusion time step $k\sim\text{Uniform}(\{1,\dots,K\})$, noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$;
       Update $\theta$ using the loss function (7);
// fine-tuning model for downstream tasks
for $\mathcal{T}_{i}\in[\mathcal{T}_{1},\dots,\mathcal{T}_{N}]$ do
       Initialization: $\theta\leftarrow\theta_{\text{PT}}$; $\mathcal{B},\mathcal{B}_{\text{target}}\leftarrow$ Rollout $N_{\text{init}}$ proficient trajectories using $\theta$;
       for t $=1,\dots,N_{\text{FT}}$ do
@@
                   Obtain samples $\bm{a}_{t}^{0:K}\sim p_{\theta}(\bm{a}_{t}^{0:K}|s_{t})$;
                   Execute the first $T_{a}$ steps and get reward $r(\bm{a}_{t}^{0})$;
                   $\mathcal{B}\leftarrow\mathcal{B}\cup(s_{t},\bm{a}_{t}^{0:K},r(\bm{a}_{t}^{0}))$;
                   $s_{t}\leftarrow s_{t+T_{a}},t\leftarrow t+T_{a}$;
            // approximate target policy $\mu$
             if proficient then
@@
            Compute $\mathcal{L}_{\text{Imp}}^{\mathcal{T}_{i}}$ using batches from $\mathcal{B}$ according to Eq. (12);
@@
             Compute $\mathcal{L}_{\text{BC}}^{\mathcal{T}_{i}}$ using batches from $\mathcal{B}_{\text{target}}$ according to Eq. (14);
             Update $\theta$ using the loss function (15);
Appendix C The Details of Baselines
@@
We describe the details of baselines used for comparison in our experiments. For Meta-World, we consider following baselines:
@@
For Adroit, we consider following baselines:
  • BCRNN. A variant of BC that models the policy network as an RNN. The network is trained on temporal sequences of length $H$, denoted as $(s_{t},a_{t},...,s_{t+H},a_{t+H})$, to predict action sequences based on historical states.
  • IBC. BC is represented as a conditional energy-based modeling problem, where implicit policies are trained to imitate expert demonstrations.
  • Diffusion Policy. The generation of robot behaviors is formulated as a conditional denoising diffusion process, where the diffusion model predicts action sequences based on given observations as conditions.
  • DP3. Diffusion Policy is extended by incorporating 3D visual representations. The 3D scenes from the environment are represented as point clouds, which are then cropped and downsampled to reduce redundant information. These processed point clouds are passed through an MLP to generate visual representations, which serve as conditions for the diffusion models.
Appendix D Variants of SODP
In Eq. (15), we introduce a BC regularization term to preserve the pre-trained knowledge and demonstrate its effectiveness compared to two existing regularization approaches presented in DPOK(Fan et al., 2024) and DLPO (Chen et al., 2024). Specifically, the regularization term $\mathcal{L}_{\text{KL}}$ used in DPOK is expressed as:
@@ Appendix D Variants of SODP
And the regularization term $\mathcal{L}_{\text{PL}}$ used in DLPO is expressed as:
@@
These methods can be considered as different approaches to selecting the target policy in line with our analysis and can be seen as variants of Eq. (14), where DPOK selects $\mu=\theta_{\text{pre-train}}$ and DLPO selects $\mu=\theta$. The rationale behind their selection is based on the assumption that $\theta\approx\theta_{\text{pre-train}}\approx\theta^{*}$. This assumption is reasonable in text-to-image or text-to-speech tasks, as the pre-trained models they used are already strong and perform exceptionally well even without fine-tuning. However, this assumption does not apply to our pre-trained planner, as the model is trained on sub-optimal data. As a result, as shown in Section 5.3, these regularization methods may lead the pre-trained planner to be stuck in inferior regions, limiting its ability to improve performance.
Appendix E Comparison to DPPO
We summarize some similarities and differences between our work and the concurrent work DPPO (Ren et al., 2024) as follows:

### v3

@@ 6 Conclusion
Acknowledgements
This work is supported by the National Key Research and Development Program of China (Grant No.2024YFE0210900), the National Natural Science Foundation of China (Grant No.62306242), the Young Elite Scientists Sponsorship Program by CAST (Grant No. 2024QNRC001), and the Yangfan Project of the Shanghai (Grant No.23YF11462200).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Derivations
A.1 Derivation of policy gradient in Equation  (3.2)
@@
A.2 Derivation of loss function in Equation (3.2)
@@ Appendix B The Details of SODP
B.1 Theoretical Motivation
The theoretical motivation of SODP is grounded in the provable efficient reward-free exploration in RL (Wang et al., 2020; Jin et al., 2020). This approach employs a two-stage learning process for developing an RL policy: a reward-free exploration phase and a reward-based optimization phase. The latter stage requires only a limited number of reward labels.
Specifically, in the reward-free stage, the agent optimizes a pure exploration reward, namely, the upper-confidence bound (UCB) term of the value function. Under the linear MDP assumption, we denote the feature of the state-action pair as $\phi(s,a)$, and the value function is a linear mapping of $\phi(s,a)$, expressed as $Q(s,a)=\phi(s,a)^{\top}\theta$, where $\theta$ is a weight vector. The covariance matrix for state-action pairs is defined as $\Lambda=\sum_{i}\phi(s_{i},a_{i})\phi(s_{i},a_{i})^{\top}$. Consequently, the UCB-based bonus for reward-free exploration is defined as $b_{t}=[\phi(s_{t},a_{t})^{\top}\Lambda^{-1}\phi(s_{t},a_{t})]^{1/2}.$ Leveraging this bonus, the agent can collect diverse state-action pairs with broad coverage during the exploration stage, thereby providing a sufficient amount of information for the subsequent planning phase. In the second stage, the agent performs least-square value iteration (LSVI), a well-studied RL algorithm, by interacting with the environment. The goal of this stage is to conduct value and policy updates for a specific task with a specific reward function. This two-stage learning paradigm, which combines reward-free exploration and reward-based optimization, can achieve polynomial sample complexity according to theoretical results (Wang et al., 2020). We believe this approach provides a solid theoretical foundation for our method’s empirical success.
Our algorithm (i.e., SODP) essentially follows this theoretical motivation but incorporates more practical implementations. Specifically, previous theoretical works have shown that $b_{t}$ defined in linear MDPs is approximately proportional to the reciprocal pseudo-count of the corresponding state-action pair in the dataset. As a result, the reward-free exploration stage in (Wang et al., 2020) is similar to count-based exploration, which aims to collect diverse datasets in the first stage (Ostrovski et al., 2017). In SODP, we adopt an alternative exploration method (entropy-based exploration) via a non-expert SAC policy to provide sufficient exploration in the data collection stage. We also pre-train a diffusion planner on this exploratory dataset rather than directly performing reward-based policy optimization. This approach seeks to model the distribution of diverse trajectories with broad coverage and provides a well-initialized starting point for the second stage. In Section 5.3, we empirically demonstrate that SODP requires only a very limited number of online samples to achieve strong performance in the second stage. This finding verifies the theoretical results of (Wang et al., 2020; Jin et al., 2020), which suggest that such a two-stage learning paradigm is sample-efficient. We believe that our practical implementation of this theoretical framework not only enhances the applicability of our method but also maintains the desirable properties of sample efficiency and robustness.
B.2 Diffusion Policy
@@ B.3 Implementation Details
  • For fine-tuning, we use DDIM (Song et al., 2021) with $10$ sampling steps and $\eta=1$. We fine-tune each task for $1e^{6}$ steps in Meta-World and $3e^{3}$ steps in Adroit. Following DPOK (Fan et al., 2024), we perform $p_{\text{step}}\in\{10,30\}$ gradient steps per episode. We set discount factor $\gamma=1$ for all tasks.
  • We set $N_{\text{init}}\in\{10,20\}$ for approximating target distribution and $\lambda=1.0$ as the BC weight coefficient.
@@
       Sample $(\bm{s},\bm{a})\sim D$, diffusion time step $k\sim\text{Uniform}(\{1,\dots,K\})$, noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$  Update $\theta$ using the loss function (7)
// fine-tuning model for downstream tasks for $\mathcal{T}_{i}\in[\mathcal{T}_{1},\dots,\mathcal{T}_{N}]$ do
       Initialization: $\theta\leftarrow\theta_{\text{PT}}$; $\mathcal{B},\mathcal{B}_{\text{target}}\leftarrow$ Rollout $N_{\text{init}}$ proficient trajectories using $\theta$for t $=1,\dots,N_{\text{FT}}$ do
@@
                   Obtain samples $\bm{a}_{t}^{0:K}\sim p_{\theta}(\bm{a}_{t}^{0:K}|s_{t})$  Execute the first $T_{a}$ steps and get reward $r(\bm{a}_{t}^{0})$$\mathcal{B}\leftarrow\mathcal{B}\cup(s_{t},\bm{a}_{t}^{0:K},r(\bm{a}_{t}^{0}))$$s_{t}\leftarrow s_{t+T_{a}},t\leftarrow t+T_{a}$
            // approximate target policy $\mu$ if proficient then
@@
            Compute $\mathcal{L}_{\text{Imp}}^{\mathcal{T}_{i}}$ using batches from $\mathcal{B}$ according to Eq. (3.2)  Compute $\mathcal{L}_{\text{BC}}^{\mathcal{T}_{i}}$ using batches from $\mathcal{B}_{\text{target}}$ according to Eq. (14)  Update $\theta$ using the loss function (15)
Appendix C Extended Results
In this section, we provide our full experimental results:
  1. 1.
    Ablation studies evaluating various fine-tuning strategies.
  2. 2.
    Analysis of the impact of pre-training dataset quality.
  3. 3.
    Additional experiments for offline-online baselines
  4. 4.
    Evaluation on image-based environments.
  5. 5.
    Computational efficiency analysis
  6. 6.
    Learning curves of SODP compared to multi-task RL baselines.
C.1 Ablation studies examining other fine-tuning approaches
To demonstrate the effectiveness of our online fine-tuning approach, we compare it with two alternative fine-tuning methods: (i) SODP_off, which involves fine-tuning using high-quality offline data, and (ii) SODP_off_scratch, which performs direct training with high-quality data without pre-training. Specifically, we fine-tuned the pre-trained models for 100k steps across five tasks, collecting 200 successful episodes (equivalent to 100k steps) for each task. These datasets were then used to independently train five models in an offline setting, utilizing the same loss function as in Eq. (15) (SODP_off). Additionally, to investigate the impact of pre-training on offline fine-tuning, we trained the model directly without pre-training (SODP_off_scratch).
The experimental results, presented in Table 4, report the success rates averaged over three seeds. Without pre-training, the model lacks the necessary action priors to efficiently identify high-reward action distributions. Furthermore, directly fine-tuning with high-quality offline data proves insufficient, as static reward labels may fail to provide adequate guidance in dynamic environments, hindering the model’s ability to facilitate efficient exploration.
Table 4: Average success rate for different fine-tuning approaches.
Tasks SODP_off SODP_off scratch SODP
button-press-topdown $58.67_{\pm 0.03}$ $40.67_{\pm 0.08}$ $60.67_{\pm 0.03}$
hammer $71.33_{\pm 0.05}$ $13.33_{\pm 0.06}$ $73.33_{\pm 0.03}$
handle-pull-side $60.67_{\pm 0.03}$ $42.67_{\pm 0.08}$ $81.67_{\pm 0.07}$
peg-insert-side $25.33_{\pm 0.03}$ $0.0_{\pm 0.0}$ $32.67_{\pm 0.06}$
handle-pull $66.67_{\pm 0.03}$ $31.33_{\pm 0.06}$ $75.33_{\pm 0.04}$
Average success rate $56.53_{\pm 0.18}$ $25.6_{\pm 0.18}$ $64.73_{\pm 0.19}$
To highlight the importance of modeling the diffusion process as a MDP for reward fine-tuning, we consider an alternative approach that directly applies BC during fine-tuning, using only Eq. (14) as the loss function. As shown in Table 5, directly using BC results in poorer performance, as BC lacks reward labels to effectively guide exploration. While BC during the fine-tuning phase enables access to dynamic actions, it is limited to ’imitation’ rather than ’evolution,’ as the model is unable to differentiate between good and bad actions.
Table 5: Average success rate for directly BC during fine-tuning.
Task Directly BC SODP
button-press-topdown $51.3_{\pm 0.05}$ $60.7_{\pm 0.03}$
basketball $21.3_{\pm 0.03}$ $41.2_{\pm 0.16}$
stick-pull $26.7_{\pm 0.08}$ $50.5_{\pm 0.04}$
C.2 Pre-training using near-optimal data
To evaluate the impact of pre-training data quality on fine-tuning performance, we modified the near-optimal dataset provided by (He et al., 2023) by retaining only the last 50% of the data. This modification ensured that the total number of transitions remained the same as the sub-optimal data used in the main paper, while significantly increasing the proportion of expert trajectories. We refer to this modified dataset as near-optimal data and pre-trained a model on the Meta-World 10 tasks. Subsequently, we followed the same fine-tuning procedure outlined in the main paper to fine-tune the model on each task. The experimental results are presented in Table 6. Incorporating more optimal data during the pre-training stage leads to better performance, as the model gains more priors about the optimal action distributions.
Table 6: Average success rate achieved after fine-tuning models pre-trained on different datasets.
Tasks Sub-optimal dataset Near-optimal dataset
basketball $52.67_{\pm 0.03}$ $80.67_{\pm 0.03}$
button-press $88.00_{\pm 0.02}$ $89.33_{\pm 0.03}$
dial-turn $80.67_{\pm 0.02}$ $74.00_{\pm 0.04}$
drawer-close $100.00_{\pm 0.00}$ $100.00_{\pm 0.00}$
peg-insert-side $62.67_{\pm 0.02}$ $84.67_{\pm 0.02}$
pick-place $36.67_{\pm 0.03}$ $59.33_{\pm 0.03}$
push $33.33_{\pm 0.03}$ $50.67_{\pm 0.03}$
reach $68.67_{\pm 0.05}$ $95.33_{\pm 0.01}$
sweep-into $60.67_{\pm 0.03}$ $75.33_{\pm 0.01}$
window-open $69.33_{\pm 0.04}$ $100.0_{\pm 0.00}$
Average success rate $65.27_{\pm 0.21}$ $80.93_{\pm 0.16}$
C.3 Additional experiments for offline-online baselines
We consider more baselines that are diffusion-based and support offline pre-training to online RL fine-tuning procedures. Specifically, we extend two diffusion-based methods, DQL (Wang et al., 2023) and IDQL (Hansen-Estruch et al., 2023), to multi-task offline-to-online training on Meta-World 10 tasks, following (Ren et al., 2025). Specifically, we first pretrain a diffusion-based actor using behavior cloning on an offline dataset and then fine-tune both the actor and critic in an online environment. The results are shown in Table 7 , the diffusion actor can be misled by inaccurate value network estimations, causing rapid forgetting of actions learned from offline pretraining. Furthermore, since goals in Meta-World are randomized, traditional methods struggle to learn how to complete them effectively.
We also consider transformer-based RL baselines. Specifically, we extend the multi-task Decision Transformer (DT) following (Zheng et al., 2022) by pre-training it on the same sub-optimal data and fine-tuning each task online for 1M steps on the Meta-World 10 tasks. We consider two types of DT: (1) the vanilla DT, which uses a deterministic policy, and (2) the variant proposed in (Zheng et al., 2022), which employs a stochastic policy. The results are presented in Table 7. The stochastic policy improves the success rate by enabling greater exploration in the online environment and discovering more valuable action patterns. However, it still underperforms compared to our method, as the diffusion model can generate more diverse actions.
Table 7: Average success rate of MT-10 tasks for diffusion-based and transformer-based offline-online baselines.
Methods Success Rate
Diffusion-based MT-DQL $10.80_{\pm 0.11}$
MT-IDQL $13.57_{\pm 0.12}$
Transformer-based MT-ODT_deter $40.07_{\pm 0.15}$
MT-ODT_stoc $50.47_{\pm 0.21}$
SODP (ours) $65.27_{\pm 0.21}$
C.4 Experiments in image-based Meta-World and Adroit
The Adroit benchmark includes three dexterous manipulation tasks, requiring a 24-degree-of-freedom dexterous hand to solve complex challenges such as in-hand manipulation and tool use. The goals in this environment are also randomized. For Adroit, we use images as the observation to assess whether our method can scale to high-dimensional input. To collect offline dataset, we train a VRL3 (Wang et al., 2022) agent for each task and use the initial 30% experiences (90K transitions) from the converged replay buffer.
The action space for different tasks in Adroit is different and is incompatible with MTDIFF and HarmoDT. Therefore, we compare SODP with following baselines designed for complex environments: (1) BCRNN(Mandlekar et al., 2021). A variant of BC that employs a Recurrent Neural Network (RNN) as the policy network, predicting the sequence of actions based on the sequence of states as input. (2) IBC(Florence et al., 2022). Extended BC with energy-based models (EBM) to train implicit behavioral cloning policies. (3) Diffusion Policy(Chi et al., 2023). A diffusion-based approach that predicts future action sequences based on historical states. (4) DP3(Ze et al., 2024). A visual imitation learning algorithm that incorporates 3D visual representations into diffusion policies, using a point clouds encoder to process visual observations into visual features. The results for these baselines are directly replicated from those reported in DP3 (Ze et al., 2024).
We scale our method to image-based observations using the Adroit benchmark by employing a point-cloud encoder from DP3 (Ze et al., 2024) to process the 3D scene represented by point clouds. Specifically, we capture depth images directly from the environment and convert them into point clouds using Open3D (Zhou et al., 2018). These point clouds are then processed by the DP3 Encoder, which maps them into visual features. We then train our diffusion planner following the same procedure in Algorithm 1 except the input states are visual features. Following DP3 (Ze et al., 2024), We compute the average of the highest 5 evaluation success rates during training and report the mean and std across 3 seeds. As shown in Table 8, our method achieves an 8.2% improvement across all tasks. Since hammer is more challenging than door, our method may need more insightful priors from pre-training to achieve better performance.
Adroit
[…]
BCRNN $0_{\pm 0}$ $0_{\pm 0}$ $9_{\pm 3}$ 3.0
BCRNN+3D $8_{\pm 14}$ $0_{\pm 0}$ $8_{\pm 1}$ 5.3
IBC $0_{\pm 0}$ $0_{\pm 0}$ $9_{\pm 2}$ 3.0
IBC+3D $0_{\pm 0}$ $0_{\pm 0}$ $10_{\pm 1}$ 3.3
Diffusion Policy $48_{\pm 17}$ $50_{\pm 5}$ $25_{\pm 4}$ 31.7
Simple DP3 $100_{\pm 0}$ $58_{\pm 4}$ $46_{\pm 5}$ 68.0
DP3 $100_{\pm 0}$ $62_{\pm 4}$ $43_{\pm 6}$ 68.3
SODP (ours) $67_{\pm 6}$ $96_{\pm 1}$ $59_{\pm 4}$ 73.9
We further conduct ablation studies on three Adroit environments, fine-tuning the same pre-trained model with and without BC regularization separately. As shown in Table 9, the performance improvement of the model fine-tuned without BC is marginal, whereas the model fine-tuned with BC achieves significantly better results. This indicates that BC also facilitates learning in high-dimensional single-task settings. This improvement can be attributed to the update stride constraint imposed by BC, which ensures that updates are not too aggressive, thereby preventing the model from becoming trapped in a suboptimal region.
Table 9: Ablation study on the image-based Adroit benchmark. We report the average success rate for the pre-trained model, as well as the results of fine-tuning the same model with and without BC regularization.
Adroit
[…]
Pre-trained model $20_{\pm 8}$ $55_{\pm 8}$ $28_{\pm 3}$ 34.3
Fine-tuned model w/o BC $38_{\pm 3}$ $63_{\pm 3}$ $30_{\pm 5}$ 43.7
Fine-tuned model w BC (ours) $67_{\pm 6}$ $96_{\pm 1}$ $59_{\pm 4}$ 73.9
Table 10: Average success rate of image-based MT-10 tasks.
Methods Success rate
DP3 $32.6_{\pm 0.23}$
MTDIFF_3D $38.0_{\pm 0.82}$
SODP $47.5_{\pm 0.18}$
We also conduct experiments on image-based Meta-World 10 tasks. Since no existing image-based sub-optimal dataset for Meta-World is available, we collect data for the 10 tasks by training separate SAC agents for each task, as done in (He et al., 2023), and rendering the environments to obtain image data. We then follow the same procedure as in Adroit to convert the images into point clouds and use the DP3 encoder to extract visual features. For comparison, we consider the following baselines: DP3 and MTDIFF_3D, an extended variant of MTDIFF that employs the same 3D visual encoder used in SODP. The experimental results are presented in Table 10. Our method also outperforms other baselines and achieves a higher success rate, demonstrating its generalizability to complex input modalities.
C.5 Computational efficiency for training and evaluating
We conduct a computational efficiency analysis comparing three offline pre-training followed by online fine-tuning baselines and our method. For training time, we report the average fine-tuning duration per task over 1 million environment steps. For evaluation time, we report the average inference time per task over 50 episodes. All results are obtained using a single NVIDIA RTX 4090 GPU.
Table 11: Computational efficiency for olline-online baselines and our method.
Methods Fine-tuning time Inference time
RLPD $4.5-5,5h$ $37s$
IBRL $6-7.5h$ $39s$
Cal-QL $8.5-10h$ $40s$
SODP (ours) $11-13h$ $47s$
C.6 Learning curves of multi-task baselines
We sample 10 tasks and present the learning curves of SODP alongside two leading baselines, MTDIFF and HarmoDT, across five random seeds. The results are shown in Figure 8, where the X-axis represents gradient steps. The first $5e^{5}$ steps (indicated by the black vertical dashed line) correspond to the pre-training phase of SODP, while the subsequent steps represent the fine-tuning phase (marked by the orange vertical dashed line).
Figure 8: Learning dynamics. We sample 10 tasks and present the learning curves of SODP, MTDIFF, and HarmoDT across five seeds. X-axis represents gradient steps. We pre-train the planner on 50M transitions, followed by fine-tuning with 1M transitions per task. SODP rapidly converges to high success rates, whereas MTDIFF and HarmoDT struggle with some challenging tasks.
SODP rapidly converges to high success rates, surpassing the other two baselines. The pre-training stage equips the planner with comprehensive action distribution priors and allows it to rapidly transfer and enhance these capabilities across a variety of downstream tasks. As a result, the pre-training stage significantly accelerates convergence, leading to more efficient learning in the fine-tuning stage. The two baseline approaches struggle to address complex and challenging tasks such as basketball and hammer. In contrast, our method effectively guides the model to generate proficient actions, demonstrating the benefits of fine-tuning with policy gradient concerning return maximization. Moreover, while HarmoDT exhibits instability across different random seeds, our method demonstrates robustness against randomness.
Appendix D The Details of Baselines
@@
We describe the details of baselines used for comparison in our experiments. For offline-online RL, we consider following baselines and use the implementation from https://github.com/irom-princeton/dppo:
  • Cal-QL. The actor network is implemented as a four-layer MLP and is pre-trained on an offline dataset using the SAC approach. During online fine-tuning, it optimizes conservative value functions that are constrained to exceed the value function of the reference policy.
  • IBRL. The actor network is implemented as a three-layer MLP and is pre-trained using imitation learning. During fine-tuning, it leverages an ensemble of $N$ critics, randomly selecting two of them for each update step.
  • RLPD. The actor network is implemented as a three-layer MLP and is pre-trained on an offline dataset using the SAC approach. During online fine-tuning, it leverages an ensemble of $N$ critics, randomly selecting two of them for each update step.
For multi-task RL, we consider following baselines:
  • MTSAC. The one-hot encoded task ID is incorporated into the original SAC as an additional input.
@@
  • MTDQL. Similar to MTIQL, a multi-head critic network is utilized to predict the Q-value for each task, and the original diffusion actor is extended with an additional task ID input.
@@
Appendix E Variants of SODP
In Eq. (15), we introduce a BC regularization term to preserve the pre-trained knowledge and demonstrate its effectiveness compared to existing regularization approaches. We detailed other regularization methods below.
  • SODP w/o regularization. This variant is similar to DDPO (Black et al., 2024) and DPPO (Ren et al., 2025), which fine-tunes the model directly using Eq. (3.2) without any regularization.
  • SODP_kl. This variant is similar to DPOK (Fan et al., 2024), with the addition of a KL regularization term to constrain the divergence between the fine-tuned model and the pre-trained model. Specifically, the regularization term $\mathcal{L}_{\text{KL}}$ used in DPOK is expressed as:
@@ Appendix E Variants of SODP
  • SODP_pl. This variant is similar to DLPO (Chen et al., 2024), incorporating the original diffusion pre-training loss (PL) into the fine-tuning objective to prevent the model from deviation. Specifically, the regularization term $\mathcal{L}_{\text{PL}}$ used in DLPO is expressed as:
@@
These methods can be considered as different approaches to selecting the target policy in line with our analysis and can be seen as variants of Eq. (14), where DPOK selects $\mu=\theta_{\text{pre-train}}$ and DLPO selects $\mu=\theta$. The rationale behind their selection is based on the assumption that $\theta\approx\theta_{\text{pre-train}}\approx\theta^{*}$. This is reasonable for methods like DPOK and DLPO, which utilize pre-trained models such as Stable Diffusion (Rombach et al., 2022) and WaveGrad2 (Chen et al., 2021b). These models already exhibit strong generative capabilities without fine-tuning, and the goal is to make slight adjustments to align them with more fine-grained attributes, such as aesthetic scores and human preferences.
However, this assumption does not apply to our pre-trained planner, as the model is pre-trained on sub-optimal data and lacks the ability to solve complex tasks. We expect it to develop new skills for completing these tasks through fine-tuning. However, directly applying KL regularization to the pre-trained model leads to conservative policies that heavily rely on the existing capability, thereby confining the model to a sub-optimal region. While PL regularization allows some slight exploration, it is uncontrolled. Consequently, as shown in Section 5.4, we observe that the KL regularization almost remains unchanged and the PL regularization slightly increases the performance in basketball but decreases in other tasks.
Appendix F Comparison to DPPO
We summarize some similarities and differences between our work and the concurrent work DPPO (Ren et al., 2025) as follows: